OpenStudy (anonymous):

1/3 (6 - 15x) = -2 + 4x Please Give me an Explanation and Answer

OpenStudy (shamil98):

1/3 ( 6 - 15x) = -2 + 4x the 1/3 is distributed to what is in the parentheses This becomes. 2 - 5x = -2x + 4x.

OpenStudy (shamil98):

2 - 5x = -2 +4x*

OpenStudy (shamil98):

From there just solve for x. Can you do that?

OpenStudy (charlotte123):

I shall be the checker ^_^ I got the answer :3

OpenStudy (charlotte123):

sham ლ(́◉◞౪◟◉‵ლ) what happened to your shiny L? ;-;

OpenStudy (anonymous):

I think. As you can see, I can't quite grasp math.

OpenStudy (charlotte123):

2 - 5x = 4x - 2 right sham? ^_^

OpenStudy (shamil98):

2 - 5x = 4x - 2 you need to get x on one side. so let's add 5x to both sides. 2 = 9x - 2 add 2 4=9x divide both sides by 9 . 4/9 = x

OpenStudy (anonymous):

Thanks bro, you're a serious help! Have a great day!

OpenStudy (charlotte123):

exactly sham ^_^

OpenStudy (charlotte123):

bye User :3

OpenStudy (anonymous):

Good-bye.

OpenStudy (anonymous):

<3

OpenStudy (charlotte123):

XD

OpenStudy (anonymous):

:p